# intersection.bib

### v2

@@ 3 Bounds for the furthest distance a car can come from
From now on we will work with $\mathcal{E}_{i}$’s. Consider the points $n_{i}^{\prime},n_{j}^{\prime}\in\mathcal{L}_{-n}$ defined as before. We have $|n_{i}^{\prime}-n_{j}^{\prime}|$ is of order $M|i-j|n^{2/3}.$ For simplicity let $k_{i,j}=M|i-j|$. We apply again an averaging argument. Consider the following parallelogram (see Figure 5).
$W:=\{w\in\mathbb{Z}^{2}:|\phi(w)|\leq\lceil\frac{n}{100k_{i,j}}\rceil,|\psi(w)|<\lceil\frac{n^{2/3}}{2}\rceil\}.$
Then for all $w\in W$ define the following events.
@@
$\frac{c_{n}n^{5/3}}{100k_{i,j}}\mathbb{P}(\mathcal{E}_{i}\cap\mathcal{E}_{j})=\sum_{w\in W}\mathbb{P}(\mathcal{E}_{i}^{w}\cap\mathcal{E}_{j}^{w})=\mathbb{E}\left(\sum_{w\in W}\mathbbm{1}_{\mathcal{E}_{i}^{w}\cap\mathcal{E}_{j}^{w}}\right),$ (3.10)
where $\{c_{n}\}$ is a uniformly bounded sequence such that for each $n,$ the number of vertices in $W$ is $\frac{c_{n}n^{5/3}}{100k_{i,j}}$. Let us define the following random variable.
  • $N_{i,j}$ is number of $w\in W$ such that there exist $u_{i},u_{j}\in\mathbb{Z}^{2}$ with $|\phi(u_{i})+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $|\phi(u_{j})+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $\theta_{u_{i}}$ (resp. $\theta_{u_{j}}$) lies in $A_{i}$ (resp. $A_{j}$) and $w\in\Gamma_{u_{i}}^{\theta_{u_{i}}}\cap\Gamma_{u_{j}}^{\theta_{u_{j}}}.$
@@
$\sum_{w\in W}\mathbbm{1}_{\mathcal{E}_{i}^{w}\cap\mathcal{E}_{j}^{w}}\leq N_{i,j}.$
We will show that there exists a constant $C>0$ (depending only on $\epsilon$) such that for all $n$ sufficiently large
@@
We define two equivalence classes. For $u,v\in\mathbb{Z}^{2}$ with $|\phi(u)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $|\phi(v)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $\theta_{u},\theta_{v}\in A_{i}$ we say $u\sim v$ if $\Gamma_{u}^{\theta_{u}}$ and $\Gamma_{v}^{\theta_{v}}$ coincide inside $W.$ Let $\widetilde{E_{i}}$ denote the number of equivalence classes. Similarly, we define the equivalence classes corresponding to $A_{j}$ and denote it by $\widetilde{E_{j}}.$ Further for $u,v\in\mathbb{Z}^{2}$ with $|\phi(u)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil,\theta_{u}$ (resp. $\theta_{v}$) contained in $A_{i}$ (resp. $A_{j}$) let $I_{u,v}$ denote the intersection size of $\Gamma_{u}^{\theta_{u}}$ and $\Gamma_{v}^{\theta_{v}}$ inside $W$. Clearly,
$N_{i,j}\leq\widetilde{E_{i}}\widetilde{E_{j}}\max I_{u,v},$
@@
where the maximum is taken over all $u,v\in\mathbb{Z}^{2}$ with $|\phi(u)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $|\phi(v)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $\theta_{u}$ (resp. $\theta_{v}$) lies in $A_{i}$ (resp. $A_{j}$). So, for $\ell\geq 1$
$\mathbb{P}\left(N_{i,j}\geq\frac{\ell n\log|k_{i,j}|}{k_{i,j}^{3}}\right)\leq\mathbb{P}(\widetilde{E_{i}}\geq\ell^{1/3})+\mathbb{P}(\widetilde{E_{j}}\geq\ell^{1/3})+\mathbb{P}\left(\max I_{u,v}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\right).$ (3.12)
@@
We start with an upper bound for the first term. Let us consider the following parallelogram (see Figure 5).
@@
  • $\mathsf{R}_{i}^{\ell}$ is the parallelogram whose opposite sides lie on $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $\mathcal{L}_{-n+\lceil\frac{n}{100k_{i,j}}\rceil}$), each with $\lceil\ell^{1/96}n^{2/3}\rceil$ many vertices with midpoints are the intersection points of $y=\tan(\theta_{i})x$ with $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $\mathcal{L}_{-n+\lceil\frac{n}{100k_{i,j}}\rceil}$).
@@
Figure 5: To find an upper bound for the second sum in (4.3) we fix $1\leq i<j\leq[\frac{n^{1/3}}{2M}]$. Due to planarity and large transversal fluctuation there is a very small probability that there will be vertices $u$ outside $\mathsf{R}_{i}^{\ell}$ (resp. $\mathsf{R}_{j}^{\ell}$) with $|\phi(v)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ choosing angle in $A_{i}$ (resp. $A_{j}$) such that $\Gamma_{u}^{\theta_{u}}$ will intersect $W$. These are the events $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$. Further, the same argument as in [BBB23, Lemma 3.1] shows that, there is a very small probability that for any geodesic starting from $\mathsf{R}_{i}^{\ell}$(resp. $\mathsf{R}_{j}^{\ell}$) and taking angle in $A_{i}$ (resp. $A_{j}$) will not intersect either $\mathcal{J}_{i}^{\prime}$ or $\mathcal{J}_{i}^{\prime\prime}$ (resp. either $\mathcal{J}_{j}^{\prime}$ or $\mathcal{J}_{j}^{\prime\prime}$). These are the events $\mathcal{G}_{i}$ and $\mathcal{G}_{j}.$ Further, due to Proposition 3.4, we have with very small probability that there are more than $\ell^{1/3}$ distinct vertices in $W$ that has a geodesic starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$). Finally, we invoke [BBB23, Lemma 3.2, Lemma 3.3] to conclude that with very small probability any geodesic starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$) will have more than $\frac{\ell n}{k_{i,j}^{3}}$ intersection size inside $W$. Combining all the above we get stretched exponential upper bound for each terms in (3.12).
@@
Let us consider the following two line segments $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{i}^{\prime\prime}$) with $\lceil 4\ell^{1/96}n^{2/3}\rceil$ many vertices on $\mathcal{L}_{-n+\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $\mathcal{L}_{n}$) with midpoints the intersection point of $y=\tan(\theta_{i})x$ and $\mathcal{L}_{-n+\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $n_{i}^{\prime\prime}$). Let us define following two events (see Figure 5).
@@
  • $\mathcal{F}_{i}:=\{\exists u\notin\mathsf{R}_{i}^{\ell}:|\phi(u)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil,\theta_{u}\in A_{i},\Gamma_{u}^{\theta_{u}}$ intersects $W\}$.
@@
  • $\mathcal{G}_{i}:=\{\exists u\in\mathsf{R}_{i}^{\ell}:\theta_{u}\in A_{i},\Gamma_{u}^{\theta_{u}}$ does not intersect either $\mathcal{J}_{i}^{\prime}$ or $\mathcal{J}_{i}^{\prime\prime}\}$.
Similarly, we define $\mathcal{F}_{j},\mathcal{G}_{j}$.
Using a planarity argument and due to large transversal fluctuation (see also the argument we used in the upper bound to bound the probability of the event $\mathcal{A}_{i}$) we have that there exist $C,c>0$ (depending only on $\epsilon$) such that for sufficiently large $\ell$ and $n$

### v3

@@ 3 Bounds for the furthest distance a car can come from
From now on we will work with $\mathcal{E}_{i}$’s. Consider the points $n_{i}^{\prime},n_{j}^{\prime}\in\mathcal{L}_{-n}$ defined as before. We have $|n_{i}^{\prime}-n_{j}^{\prime}|$ is of order $M|i-j|n^{2/3}.$ For simplicity let $k_{i,j}=M|i-j|$. We apply again an averaging argument. We fix $i$ and $j$. Same as before we define $\theta_{i}$ and $\theta_{j}$ to be the midpoints of the intervals $A_{i}$ and $A_{j}$. We define parallelograms $W_{i,j}$ depending on $\theta_{i}$ and $\theta_{j}$ as follows. Let both $\theta_{i}>\frac{\pi}{4}$ and $\theta_{j}>\frac{\pi}{4}$ and $\theta_{i}>\theta_{j}$. Then consider the parallelogram $W_{i,j}$ whose two sides lie on the lines $\mathcal{L}_{-\lceil\frac{n}{100k_{i,j}}\rceil}$ and $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$ each of length $n^{2/3}$ with midpoints $v_{1}^{\prime},v_{2}^{\prime}$ where $v_{1}^{\prime}$ and $v_{2}^{\prime}$ are the intersection points of the line $y=\tan(\theta_{j})x$ with $\mathcal{L}_{-\lceil\frac{n}{100k_{i,j}}\rceil}$ and $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$ respectively. Now if $\theta_{i}<\frac{\pi}{4}$ and $\theta_{j}<\frac{\pi}{4}$ and $\theta_{i}>\theta_{j}$ then define the parallelogram $W_{i,j}$ exactly same as above with the only difference that the midpoints $v_{1}^{\prime},v_{2}^{\prime}$ are the intersection points of the line $y=\tan(\theta_{i})x$ with $\mathcal{L}_{-\lceil\frac{n}{100k_{i,j}}\rceil}$ and $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$ respectively. Finally, if $\theta_{i}<\frac{\pi}{4}<\theta_{j}$ then define $W_{i,j}$ exactly same as above with the midpoints $v_{1}^{\prime},v_{2}^{\prime}$ are the intersection points of the line $y=x$ with $\mathcal{L}_{-\lceil\frac{n}{100k_{i,j}}\rceil}$ and $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$ respectively (see Figures 5, 6). Then for all $w\in W_{i,j}$ define the following events.
@@
$\frac{cn^{5/3}}{100k_{i,j}}\mathbb{P}(\mathcal{E}_{i}\cap\mathcal{E}_{j})\leq\sum_{w\in W}\mathbb{P}(\mathcal{E}_{i}^{w}\cap\mathcal{E}_{j}^{w})=\mathbb{E}\left(\sum_{w\in W}\mathbbm{1}_{\mathcal{E}_{i}^{w}\cap\mathcal{E}_{j}^{w}}\right),$ (3.7)
where the number of vertices in $W_{i,j}$ is at least $\frac{cn^{5/3}}{100k_{i,j}}$ for all large $n$. Let us define the following random variable.
  • $N_{i,j}$ is number of $w\in W_{i,j}$ such that there exist $u_{i},u_{j}\in\mathbb{Z}^{2}$ and $\gamma_{i}\in A_{i}$ and $\gamma_{j}\in A_{j}$ with $|\phi(u_{i})+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $|\phi(u_{j})+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $w\in\Gamma_{u_{i}}^{\gamma_{i}}\cap\Gamma_{u_{j}}^{\gamma_{j}}.$
@@
$\sum_{w\in W_{i,j}}\mathbbm{1}_{\mathcal{E}_{i}^{w}\cap\mathcal{E}_{j}^{w}}\leq N_{i,j}.$
We will show that there exists a constant $C>0$ (depending only on $\varepsilon$) such that for all $n$ sufficiently large
@@
We define two equivalence classes. For $u_{1},u_{2}\in\mathbb{Z}^{2}$ and $\gamma^{1}_{i},\gamma_{i}^{2}\in A_{i}$ with $|\phi(u_{1})+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $|\phi(u_{2})+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ we say $(u_{1},\gamma^{1}_{i})\sim(u_{2},\gamma^{2}_{i})$ if $\Gamma_{u_{1}}^{\gamma^{1}_{i}}$ and $\Gamma_{u_{2}}^{\gamma^{2}_{i}}$ coincide inside $W_{i,j}.$ Let $\widetilde{E_{i}}$ denote the number of equivalence classes. Similarly, we define the equivalence classes corresponding to $A_{j}$ and denote it by $\widetilde{E_{j}}.$ Further for $u,v\in\mathbb{Z}^{2}$ with $|\phi(u)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil,|\phi(v)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil,\gamma_{i}$ (resp. $\gamma_{j}$) contained in $A_{i}$ (resp. $A_{j}$) let $I_{u,v,\gamma_{,}\gamma_{j}}$ denote the intersection size of $\Gamma_{u}^{\gamma_{i}}$ and $\Gamma_{v}^{\gamma_{j}}$ inside $W_{i,j}$. Clearly,
$N_{i,j}\leq\widetilde{E_{i}}\widetilde{E_{j}}\max I_{u,v,\gamma_{i},\gamma_{j}},$
@@
where the maximum is taken over all $u,v\in\mathbb{Z}^{2}$ with $|\phi(u)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $|\phi(v)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ and $\gamma_{i}$ (resp. $\gamma_{j}$) lies in $A_{i}$ (resp. $A_{j}$). So, for $\ell\geq 1$
$\mathbb{P}\left(N_{i,j}\geq\frac{\ell n\log|k_{i,j}|}{k_{i,j}^{3}}\right)\leq\mathbb{P}(\widetilde{E_{i}}\geq\ell^{1/3})+\mathbb{P}(\widetilde{E_{j}}\geq\ell^{1/3})+\mathbb{P}\left(\max I_{u,v,\gamma_{i},\gamma_{j}}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\right).$
@@
We start with an upper bound for the first term. Let us consider the following parallelogram (see Figures 5, 6).
@@
  • $\mathsf{R}_{i}^{\ell}$ is the parallelogram whose opposite sides lie on $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $\mathcal{L}_{-n+\lceil\frac{n}{100k_{i,j}}\rceil}$), each of length $\ell^{1/96}n^{2/3}$ with midpoints being the intersection points of $y=\tan(\theta_{i})x$ with $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $\mathcal{L}_{-n+\lceil\frac{n}{100k_{i,j}}\rceil}$).
Similarly we define $\mathsf{R}_{j}^{\ell}$ by replacing $\theta_{i}$ with $\theta_{j}.$
@@
Figure 5: To find an upper bound for the second sum in (4.3) we fix $1\leq i<j\leq[\frac{n^{1/3}}{2M}]$. Due to planarity and large transversal fluctuation there is a very small probability that there will be vertices $u$ (resp. $v$) outside $\mathsf{R}_{i}^{\ell}$ (resp. $\mathsf{R}_{j}^{\ell}$) with $|\phi(u)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ (resp. $|\phi(v)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$) and $\gamma_{i}\in A_{i}$ (resp. $\gamma_{j}\in A_{j}$) such that $\Gamma_{u}^{\gamma_{i}}$ (resp. $\Gamma_{v}^{\gamma_{j}}$) will intersect $W_{i,j}$. These are the events $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$. Further, the same argument as in [BBB23, Lemma 3.1] shows that, there is a very small probability that for any geodesic starting from $\mathsf{R}_{i}^{\ell}$(resp. $\mathsf{R}_{j}^{\ell}$) and taking angle in $A_{i}$ (resp. $A_{j}$) will not intersect either $\mathcal{J}_{i}^{\prime}$ or $\mathcal{J}_{i}^{\prime\prime}$ (resp. either $\mathcal{J}_{j}^{\prime}$ or $\mathcal{J}_{j}^{\prime\prime}$). These are the events $\mathcal{G}_{i}$ and $\mathcal{G}_{j}.$ Further, due to Proposition 6.4, we have with very small probability that there are more than $\ell^{1/3}$ distinct vertices in $W_{i,j}$ that have a geodesic starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$). Finally, we invoke [BBB23, Lemma 3.2, Lemma 3.3] to conclude that with exponentially small probability any geodesic starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$) will have more than $\frac{\ell n}{k_{i,j}^{3}}$ intersection size inside $W_{i,j}$. Combining all the above we get stretched exponential upper bound for each terms in (3.2).
@@
Let us consider the following two line segments $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{i}^{\prime\prime}$) with $4\ell^{1/96}n^{2/3}$ many vertices on $\mathcal{L}_{-n+\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $\mathcal{L}_{n}$) with midpoints the intersection point of $y=\tan(\theta_{i})x$ and $\mathcal{L}_{-n+\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $n_{i}^{\prime\prime}$). Let us define following two events (see Figures 5, 6).
@@
  • $\mathcal{F}_{i}:=\{\exists u\notin\mathsf{R}_{i}^{\ell}\text{ and }\exists\gamma_{i}\in A_{i}:|\phi(u)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil,\Gamma_{u}^{\gamma_{i}}$ intersects $W_{i,j}\}$.
@@
  • $\mathcal{G}_{i}:=\{\exists u\in\mathsf{R}_{i}^{\ell}\text{ and }\gamma_{i}\in A_{i}:\Gamma_{u}^{\gamma_{i}}$ does not intersect either $\mathcal{J}_{i}^{\prime}$ or $\mathcal{J}_{i}^{\prime\prime}\}$.
Similarly, we define $\mathcal{F}_{j},\mathcal{G}_{j}$ with $\theta_{j}$ in place of $\theta_{i}.$ We have the following lemma.
Lemma 3.9.
There exist $C,c>0$ (depending only on $\varepsilon$) such that for sufficiently large $\ell$ and $n$